A certain mas of hydrogen gas collected over water at 6°c and 765mmHg pressure has a volume of 35cm^3, calculate the volume when its dry at STP (S.V.P of water 6°c =7mmHg)



.

1
Expert's answer
2022-03-12T01:57:27-0500

Ptotal=PH2 + PH2O

765= PH2O +7

PH2O = 758

P1V1÷T1 = P2V2÷T2

T1= 6+273= 279K

P1= 758mmHg

V1= 35cm3

T2= 273K

P2= 760mmHg

V2=?

758×35÷279= 760×V2 ÷273

V2=(758×35×273)÷(279×760)

=34.16cm3